OpenStudy (technodynamic):

f(x)=x^2-6 f(1/2) Show the work please

OpenStudy (anonymous):

\[f(\frac{1}{2}) = (\frac{1}{2})^{2} - 6\] \[f(\frac{1}{2}) = \frac{1^{2}}{2^{2}} - 6\] \[f(\frac{1}{2}) = \frac{1}{4} - \frac{6}{1}\] \[f(\frac{1}{2}) = \frac{1}{4} - \frac{24}{4}\]You have to get a common denominator to subtract the two. So finally you get \[f(\frac{1}{2}) = - \frac{23}{4}\]